OpenStudy (anonymous):

Find the equation of a line which passes through point (4, 5) and is parallel to the line formed by the equation y = -4x + 4. Express it in slope-intercept form.

OpenStudy (lalaly):

parallel lines hav the same slope y-5=-4(x-4) y-5 = -4x+16 y=-4x +21

OpenStudy (anonymous):

slope of your line is -4 use point slope formula \[y-y_1=m(x-x_1)\] withh \[m=-4,x_1=4,y_1=5\]

OpenStudy (anonymous):

and you get lalaly answer